### v1

# One Stone, Four Birds: A Comprehensive Solution for QA System Using Supervised Contrastive Learning

Abstract
This paper presents a novel and comprehensive solution to enhance both the robustness and efficiency of question answering (QA) systems through supervised contrastive learning (SCL). Training a high-performance QA system has become straightforward with pre-trained language models, requiring only a small amount of data and simple fine-tuning. However, despite recent advances, existing QA systems still exhibit significant deficiencies in functionality and training efficiency. We address the functionality issue by defining four key tasks: user input intent classification, out-of-domain input detection, new intent discovery, and continual learning. We then leverage a unified SCL-based representation learning method to efficiently build an intra-class compact and inter-class scattered feature space, facilitating both known intent classification and unknown intent detection and discovery. Consequently, with minimal additional tuning on downstream tasks, our approach significantly improves model efficiency and achieves new state-of-the-art performance across all tasks.
Index Terms:
Intent discovery, natural language processing, out-of-domain detection, question-answering system, supervised contrastive learning, text classification.
[…]
=-21pt
I Introduction
Figure 1: Our proposed workflow to achieve a comprehensive and adaptive QA system. First, an initial model is trained using current known data (Pre). Then, during daily usage, an inquiry will first be judged to be an answerable input or not (T-2): if so, proceed to the intent detection (T-1) and retrieve the answer; if not, the unknown inquiry will be temporally saved and be clustered later (T-3). Finally, the initial model is re-trained with newly discovered data to obtain the classification ability to new intents (T-4).
Question answering (QA) system is a well-established and important application widely used in areas such as healthcare [1], finance [2] and e-commerce [3]. Despite the world gradually moving into the era of large language models (LLMs), classic QA systems still maintain a strong presence due to their low cost, ease of deployment, and high accuracy in answering questions.
A typical QA system mainly comprises two components: user intent classification and answer retrieval. The efficiency of the former is crucial, as a correct response often depends on accurately determining the user’s intent. To achieve robust classification, it is essential to efficiently extract features from user input, providing intent-specific and discriminative text representations for classification. Initially, text features were extracted manually using techniques like term frequency-inverse document frequency (TF-IDF [4]) or bag-of-words (BoW [5]). With the advent of neural networks, feature extraction became an automated process [6, 7]. Today, the power of pre-trained Transformers like BERT [8] has made it much easier to extract features and build classification models, requiring only a small amount of data with Cross-Entropy (CE) loss fine-tuning [9, 10].
It seems that the current QA system building is simple and mature, with nearly nothing to be changed. However, from the perspective of building a comprehensive system, current QA systems still face unresolved issues at both internal and external levels, which have not been fully addressed yet:
  1. $\bullet$
    Externally, or at the system level: Current QA systems typically adopt a “training-as-complete” policy, meaning that once the model has finished training, its knowledge is not updated. This approach leaves the system vulnerable to unseen, unknown inputs (as out-of-domain, OOD inputs [11]). To maintain the robustness and security of the model, there should be mechanisms to handle these inputs. Furthermore, even when new data is collected to upgrade the system, the common practice is to re-train a new model from scratch, leading to a waste of labor and resources.
  2. $\bullet$
    Internally, or in training efficiency: Increasing studies indicate that the BERT+CE training combination still has efficiency problems. For BERT, the raw text feature distribution is anisotropic [12]. Additionally, when using CE as the fine-tuning loss, text features of different intents often remain mixed with unclear boundaries in the embedding space [13]. Both issues hinder training efficiency and model performance.
In this paper, aiming to build a more robust QA system, we propose a novel and comprehensive solution to address both issues. To tackle the systematic issue, including the basic intent classification function, we define four key tasks and design a workflow between them (as shown in Figure 1), enabling the system to be self-consistent and recurrently updatable. Specifically, the tasks are:
  1. T-1
    User input intent classification: The fundamental task for a QA system.
  2. T-2
    OOD input detection: Detect abnormal inputs to prevent misleading output. The detected OOD texts are saved as source data for Task 3.
  3. T-3
    New intent discovery: Utilize detected OOD texts to discover new possible intents once they accumulate to a sufficient quantity.
  4. T-4
    Continual learning: Use newly discovered and pseudo-labeled data to incrementally train the initial classification model, enhancing the system’s capability to handle the latest inquiries.
To address the training efficiency problem, we leverage supervised contrastive learning (SCL, [14]) to replace CE fine-tuning as the unified text representation optimization method, which requires minimal tuning to achieve all four task objectives. We choose SCL because, in our previous research on OOD detection [15], it demonstrated excellent capability in upstream feature optimization, allowing an accurate OOD text detection using only a simple distance-based algorithm in downstream.
Figure 2: Diagrams of SCL’s representation optimization procedure, and solutions for each task. Left: the training and optimization effect by SCL, which gathers same-intent text features together and pushes different features further apart, providing an underlying embedding space for following tasks. Right: the classification/detection processes of T-1 to T-4 after SCL. (Together with Figure 1 for better understanding)
More specifically, the advantage of SCL over CE fine-tuning lies in its ability to enhance both learning efficiency and feature separation. As shown in Figure 2-b), SCL employs a simple but clear optimization target: pulling same-label text features together and pushing different ones further apart, resulting in an intra-class compact and inter-class scattered, well-organized embedding space. This clear separation between clusters facilitates easier in-domain (IND) intent classification and IND/OOD separation and detection (Figure2, T-1/2). Additionally, this structure also supports the aggregation of currently unknown but semantically similar features, creating a conducive environment for new intent discovery and continual learning (Figure 2, T-3/4). Consequently, even with the simplest algorithms in downstream tuning, all task objectives can be achieved effortlessly and efficiently.
Our contributions can be summarized as follows:
  1. 1.
    For the first time, we clearly identify the systematic and efficiency issues in current QA systems and propose a comprehensive solution involving four key tasks in a natural workflow to enhance the training process, resulting in a robust and comprehensive system.
  2. 2.
    To achieve the task objectives, we leverage SCL as the unified representation learning approach, which requires minimal tuning in downstream tasks.
  3. 3.
    Our experiments demonstrate that, despite its conciseness and simplicity, the proposed solution significantly improves global feature optimization efficiency and outperforms state-of-the-art methods across all tasks, complementing the deficiencies of current QA systems.
II Related Works
II-A Representation Learning in Intent Classification Model Building
Early history of representation learning
Representation learning (RepL), or feature learning, is a crucial process in building machine learning models. Good RepL methods capture the intrinsic patterns and structures of the data, thus providing a solid foundation for downstream tasks.
Initially, RepL research focused on manually extracting meaningful features from raw data, with preliminary methods including TF-IDF [4] and BoW [5]. Subsequently, the development of word embeddings extracted via neural networks automated the feature extraction process [6, 7]. More recently, with the advent of large-scale encoders like BERT [8], pre-trained on extensive corpora, obtaining high-quality text embeddings has become much easier. The most common approach is to use BERT as the base encoder, with CE fine-tuning as the RepL method to refine the text embedding distribution.
Drawbacks of BERT+CE mode
The above BERT+CE approach is also widely adopted in QA systems [16, 10, 1, 17]. However, this method can be suboptimal for feature optimization from two main reasons. Firstly, from a global perspective, [18, 12] reported that the overall text embedding distribution generated by BERT is anisotropic. This means that most embeddings are spatially distributed in an elongated cone-like space rather than uniformly in a sphere, leading to high similarity between any two sentence representations and posing challenges for downstream tasks. Secondly, from a local perspective, CE fine-tuning results in a “loose” feature arrangement [13, 14], where the boundaries between different intent clusters are barely distinguishable. This not only reduces the performance of known, IND classification, but also makes the model prone to confusing IND inputs with unseen but similar OOD inputs.
Contrastive learning as a RepL method
As discussed above, RepL research has gradually shifted focus from feature extraction to feature optimization, with one of the most popular method series being the contrastive learning (ConL). The earliest ConL can be traced back to supervised metric learning (MeL) [19] used in image classification tasks, which established the fundamental strategy of “gathering similar examples together and pushing dissimilar ones further apart”. This strategy helps generate compact feature clusters efficiently. Subsequently, unsupervised ConL (UConL) [20, 21] was proposed, sharing the same strategy but aiming to improve the overall representation uniformity. Finally, SCL [14] incorporates the advantages of both MeL and UConL, thus having the potential to address both global and local feature optimization issues.
However, ConL has been mainly examined in computer vision (CV) field. In NLP field, there have been limited applications, primarily in natural language inference (NLI) tasks using UConL methods [22, 23]. A boarder exploration of SCL’s learning effect across various tasks, including its application in QA systems, remains unaddressed.
II-B Methods Review for Specific Tasks
OOD detection: As QA systems have evolved significantly in recent years, there is now a higher expectation for systems not only to answer known questions correctly but also to identify and reject unknown inputs, which constitutes the OOD detection task. Early OOD detection methods relied on the availability of numerous OOD data for binary classification [24]. However, the scope and variety of OOD data are theoretically infinite, making collection a burdensome task. Subsequently, more approaches have focused on training networks using only IND data to generate heterogeneous representations [25, 26]. Various downstream algorithms, such as local outlier factor (LOF) [27], Gaussian discriminant analysis (GDA) [28], energy score [29, 30], and Mahalanobis distance [31, 32], have been applied to distinguish between IND and OOD text features. The latest advances in OOD detection fall into two broad categories: using ConL methods to optimize the representation space [33, 34, 15], and employing generative modeling approaches like training a GPT model [35] on IND data to achieve lower likelihoods for OOD text generation [36, 37, 38].
New intent discovery, NID: Initially, new intent discovery focused on identifying new categories directly from unlabeled data, with most methods applying unsupervised clustering [39, 40]. Later approaches commonly incorporated some labeled data to support the discovery of unknown intents [41, 42]. A typical workflow involves computing textual similarities between unlabeled and known data, assigning pseudo labels to unknown data based on the results, and then performing clustering. Recently, NID research has also emphasized the importance of representation learning, with many studies incorporating pre-training and contrastive learning methods to optimize feature space arrangements [43, 44, 45].
Continual learning: Continual learning denotes a re-training procedure to let the network learn new knowledge in stages, and finally reach a level to handle all tasks. For example, we may let the model learn the even five handwriting numbers and then the odd fives, finally let it recognize all ten numbers. The biggest challenge of continual learning is the forgetting problem [46], that is, the model’s ability to recognize old data may deteriorate significantly after learning new data. Current studies are also focusing on solving this issue, where the solutions mainly include weight regularization [47, 48] and knowledge distillation [49, 50]. In this paper, for simplicity, we apply a very straightforward method that re-trains the model with both a few new and old intent data, which can be regarded as the “replay” policy [51, 52], to mitigate forgetting. Experiments (Section VI-E) show that this policy is effective and exhibits minimal performance degradation for old data classification.
Despite the above significant progress in each task, there remains a gap in integrating these tasks into a comprehensive QA system building, which is the aim of this paper to realize using SCL.
III Problem and Task Definition
As mentioned in Section I, we introduce four basic tasks to build a comprehensive and adaptive QA system (Figure 1). Here, before presenting the specific solutions, we formally define the problems and targets of each task.
III-A Preliminary
[…]
Encoder: We have a language encoder ENC that encodes and transforms a discrete text $\bm{x}$ into an m-dimension continuous representation: $\bm{x}$$\mathbb{R}^{m}$. Since there are many text similarity calculations during SCL training, we adopt a simple but effective “mean-pooling” strategy to obtain the sentence representation. Specifically, the average of all word embeddings in the sentence is served as the sentence embedding:
$\small\bm{h}=\frac{1}{len(\bm{x})}\sum_{i=1}^{len(\bm{x})}ENC(w_{i})$ (1)
where $len(\bm{x})$ is the length of input $\bm{x}$, $\{w_{1},...,w_{i},...,w_{len(\bm{x})}\}$ are the words in the sentence $\bm{x}$.
III-B T-1: User intent classification
The target of this task is to find an accurate mapping:
$f_{cls}:ENC(\bm{x}_{i})\mapsto\hat{y}_{i}\qquad\hat{y}_{i}\in\mathcal{C}_{kn}$ (2)
For traditional methods like CE-fine-tuning, this is achieved by further adding a fully-connected layer to the end of $ENC$ and training them together using CE loss. In the proposed method, the encoder is pre-trained using SCL (refer to Section IV-A for details) and the parameters are fixed in the following tasks T-1/2/3. $f_{cls}$ is realized by a nearest neighbor algorithm (Section IV-B1).
III-C T-2: Out-of-Domain detection
The target of this task is to determine whether an example $\bm{x}_{j}$ from $\mathcal{S}_{unlabeled}$ belongs to $\mathbb{D}_{IND}$ or $\mathbb{D}_{OOD}$. This is usually realized by comparing the detection score $sco$ given by a detection algorithm $f_{det}$ with an empirically-set threshold $\lambda$:
$sco_{j}=f_{det}(ENC(\bm{x}_{j}))$ (3)
$\bm{x}_{j}\in\left\{\begin{array}[]{l}\mathbb{D}_{OOD},\;if\;sco_{j}\geq\lambda\\ \mathbb{D}_{IND},\;if\;sco_{j}<\lambda\end{array}\right.$ (4)
Here, the detected $\bm{x}_{j^{\prime}}^{ood}$ are saved for T-3 (new intent discovery), while the detected $\bm{x}_{j^{\prime}}^{ind}$ will be classified normally in T-1 and also saved for the final task, T-4 (continual learning).
III-D T-3: New intent discovery
Once the collected OOD texts $\bm{x}_{j^{\prime}}^{ood}$ reach a sufficient quantity, an algorithm $f_{nid}$ is used to cluster and assign pseudo labels $\hat{y}_{j^{\prime}}$ to them, as the “new IND intents” in $\mathcal{C}_{kn^{\prime}}$:
$f_{nid}:ENC(\bm{x}_{j^{\prime}}^{ood})\mapsto\hat{y}_{j^{\prime}}\qquad\hat{y}_{j^{\prime}}\in\mathcal{C}_{kn^{\prime}}$ (5)
III-E T-4: Continual learning
[…]
$f_{cls\text{-}new}:ENC_{new}(\bm{x}_{k})\mapsto\hat{y}_{k}\qquad\hat{y}_{k}\in\{\mathcal{C}_{kn}+\mathcal{C}_{kn^{\prime}}\}$ (6)
IV Methodology
In this section, we introduce the methodology of SCL as the RepL method in the pre-training stage, followed by the downstream solutions to each task.
IV-A Stage 1: Supervised contrastive learning -based pre-training
IV-A1 Principle and optimizing effect of SCL
As discussed in Section II, BERT+CE fine-tuning exhibits issues with local loose borders and a global anisotropic distribution in RepL, while SCL has the potential to solve both of these problems.
Specifically, as illustrated in Figure 2, for the local aspect, compared with CE fine-tuning, SCL applies a more straightforward yet effective target: pulling the representations of similar examples (positive examples) together and pushing dissimilar ones (negative examples) further apart. To tackle the global distribution problem, SCL further generates “view” augmentations as additional positive examples before the contrasting process, optimizing features in a boarder scale. Together, these strategies provide a better supervision signal to arrange text representations to form compact and distinct clusters, enhancing both the alignment and uniformity of feature space, and thus sharpening the distinctions between texts from various intents.
This well-organized embedding space (Figure 2-b) provides a strong foundation that benefits all downstream tasks:
  1. T-1:
    The clear separations between each intent cluster facilitate easy and accurate classification.
  2. T-2:
    The compact and distinct IND clusters highlight the differences between them and OOD texts, which do not belong to any known intents.
  3. T-3:
    Although not trained directly with OOD data, the two optimization directions of SCL still drive the aggregation of semantically similar features. This creates an environment conducive to intent discovery algorithms, making it easier to identify the clustered gathered data.
  4. T-4:
    The accurately discovered new intent data, pre-organized feature space, and SCL-based re-training, together ensure the high-performance of the updated system.
IV-A2 Generation of text “views”
As discussed in the last subsection, the effective optimization of SCL is attributed to both the supervised positive-negative movement and the unsupervised “view” augmentation. This augmentation also serves as a precaution against the lack of positive examples in the mini-batch, especially when the data quantity is small compared with category numbers.
For images, view generation can be achieved by simple transformations such as rotating and cropping. However, these techniques are not easily applicable to text “view” generation, as hasty word swaps or deletions can distort the original meaning. In our previous study [15], we found that encoding a sentence twice with different drop-out masks (Figure 2-left) is an effective way to generate high-quality, slightly different but main meaning preserved text embeddings. By using this generation method, SCL gave the best OOD detection results compared to no-view-generation [33] or more complex techniques like embedding-attack [53] (refer to Tables 3-5 of [15] for detailed results). Considering these advantages, we continue to use this technique as the augmentation method in this paper.
IV-A3 Formal learning procedure definition
Here we give the formal definition of an SCL procedure:
Step 1: Text “View” generation
[…]
$\bm{h}_{2i^{\prime}}=ENC_{drop1}(\bm{x}_{i}),\quad\bm{h}_{2i^{\prime}+1}=ENC_{drop2}(\bm{x}_{i})$ (7)
[…]
Step 2: Supervised contrastive learning
The SCL training loss is defined as follows:
$\small{\mathcal{L}}_{SCL}\text{=}\sum_{i^{\prime}\text{=}1}^{2N}\frac{-1}{\left|Pos(i^{\prime})\right|}\sum_{p\in Pos({i^{\prime}})}\log\frac{exp(\bm{h}_{i^{\prime}}\cdot\bm{h}_{p}/\tau)}{{\displaystyle\sum_{q\in Neg({i^{\prime}})}}exp(\bm{h}_{i}^{\prime}\cdot\bm{h}_{q}/\tau)}$ (8)
[…]
IV-B Stage 2: Solutions for Downstream Tasks
With a well-organized embedding space achieved through SCL, we can complete all the tasks with the least efforts.
IV-B1 Solution for T-1 intent classification
We apply a Mahalanobis distance (MDist) - nearest neighbor (NN)-based method for IND intent classification, which does not require any additional training. We use MDist because, compared to Euclidean distance, MDist considers the distribution of each feature component, providing a more accurate measurement of text similarity. Additionally, to further accelerate the classification but avoid the instability of NN, we propose calculating MDist between the current test sample $\bm{x}_{test}$ and only the centroids $\bar{\bm{x}}_{c}$ of each intent $c$. Finally, $y_{test}$ is determined as the label of the nearest centroid. The whole process is divided into three steps:
a) Class centroid calculation
The class centroid is the average data point of all text embeddings belonging to this intent:
$\small\bar{\bm{h}}_{c}=\frac{1}{num(c)}\sum_{i=1}^{num(c)}\bm{h}_{c,i}$ (9)
where $num(c)$ is data quantity in category c, and $\bm{h}_{c,i}$ is the text representation of its i-th example.
b) MDist calculation
We calculate the covariance matrix $S$ over all data (eq. 10) and then calculate MDist (eq. 11).
$\small S=\frac{1}{\left|C\right|}\sum_{c\in{\mathcal{C}}}(\bm{H}_{c}-\overline{\bm{H}}_{c}){(\bm{H}_{c}-\overline{\bm{H}}_{c})}^{T}$ (10)
$MDist(\bm{h}_{test},\bar{\bm{h}}_{c})=\sqrt{{(\bm{h}_{test}-\bar{\bm{h}}_{c})}^{T}S^{-1}(\bm{h}_{test}-\bar{\bm{h}}_{c})}$ (11)
where $\bm{H}_{c}\text{=}[\bm{h}_{c,1},\bm{h}_{c,2},...,\bm{h}_{c,num(c)}]^{T}$ is the sample matrix of category $c$, $\overline{\bm{H}}_{c}\text{=}[1,1,..,1]^{T}\bar{\bm{h}}_{c}$.
c) Label determination
Finally, the label is determined by the nearest centroid:
$\small y_{\text{test}}=\arg\min_{c\in\mathcal{C}}MDist(\bm{h}_{\text{test}},\bar{\bm{h}}_{c})$ (12)
IV-B2 Solution for T-2 out-of-domain detection
The solution for T-2 is very simple: we still calculate MDist of $\bm{x}_{test}$ to the nearest intent centroid as T-1, but the distance is used directly as the detection $sco$:
$sco(\bm{x}_{test})={\underset{c\in{\mathcal{C}}}{min}}\ MDist(\bm{h}_{\text{test}},\bar{\bm{h}}_{c})$ (13)
The distance itself can serve as the $sco$ and be an effective detection indicator, since a larger value means that even taking the shortest distance, current example is still far from / has less similarity to any known, IND category centroids. As shown in Figure 2-T-2, this suggests that the current $\bm{x}_{test}$ is more likely to be an unseen, i.e., OOD text.
In practical usage, as introduced in Section III-C, $sco$ is further compared with a threshold $\lambda$, and texts with an exceeding distance will be judged as OOD examples.
IV-B3 Solution for T-3 new intent discovery
To conduct intent discovery, the first step is to obtain the OOD examples to be clustered. Here, we first find a threshold $\lambda$ with a policy that when the model in T-2 gets a true positive rate, TPR=90% in T-2 on validation data (i.e., 90% OOD examples are correctly recognized), then we filter out OOD texts by comparing $sco$ with this $\lambda$.
Next, since SCL has provided a well-structured feature space, we apply KMeans, one of the simplest unsupervised clustering algorithms, to solve T-3. Taking the detected OOD samples $\bm{x}_{j^{\prime}}^{ood}$, KMeans clusters and assigns pseudo-labels $\hat{y}_{j^{\prime}}$ to them. The detailed procedure in algorithm is provided in Appendix A. The pseudo-labeled new intent examples $\{\bm{x}_{j^{\prime}}^{ood},\hat{y}_{j^{\prime}}\}$ will serve as the training data for the final T-4.
IV-B4 Solution for T-4 continual learning
[…]
V Experiment
V-A Baselines
We apply the following latest and competitive methods as baselines for evaluation.
Fine-tuning, FT (All tasks): Despite being the most common training method, CE-based fine-tuning remains one of the most powerful methods actively used among all tasks [16, 31, 41]. Thus this time we applied it as a baseline of RepL, and tested the effects throughout all four tasks. The downstream detection and clustering method are all the same as SCL.
PTO-Sup [38] (T-2): This work proposed a Prefix Tuning-based OOD detection method, training a generative GPT-2 model on IND data and then using the likelihood as $sco$. It assumes that the unseen OOD text should let model output a lower likelihood. This method originally supports unsupervised learning, but for a fair comparison, we applied its supervised mode with IND labels available.
Layer-LOF [27] (T-2): This paper addressed that pre-trained language models like BERT may over-rationalize the input. That is, even it is an OOD input, the last-layer embedding may tend to be more like IND. Based on this observation, it proposed to consider a decision of multiple layers after fine-tuning, and finally used local outlier factor (LOF) with a voting mechanism for final decision. In this paper, we simply averaged its best-four layer results as the final result.
MTP-KNNCL [44] (T-3): The state-of-the-art NID method that first pre-trains the model with a Masked Language Modeling (MLM) target to obtain prior semantic knowledge, then performs KNN contrastive learning for better cluster aggregation. The downstream clustering method is KMeans.
DPN [45] (T-3): This Decoupled Prototypical Network first initializes some prototypes for known and unknown intents and aligns them as a bipartite matching problem. Unaligned prototypes represent new intents, and all embeddings are further optimized using semantic-aware prototypical learning. The downstream method is also KMeans.
V-B Datasets
We evaluate the methods using three popular intent classification datasets. To measure the performance of T-3 and T-4 quantitatively, we randomly select approximately 75% of categories as the known IND intents, treating the remaining 25% as the unknown, OOD intents to discover new intents from them. Additionally, due to a shortage of test data for T-4, we split around 20% of the training data to create a second test set, Test II.
The original data composition is shown in Table I, with examples from each dataset provided in Table V. Details on IND/OOD intent segmentation and reorganized data composition are shown in Tables II and III, respectively. To prevent confusion due to multiple data partitions and evaluation procedures, a quick reference for data usage is provided in Table IV, with a more detailed explanation in following Section V-D.
CLINC [54]: This is a widely-used dataset in OOD detection. It originally provides 150 intents of texts as the IND data and one OOD intent data (consisting of various abnormal inputs without specific labels). In this experiment, since we cannot calculate clustering metrics on this unlabeled OOD data, we discard it, and randomly select 112 IND intents as the new known IND intents and the remaining 38 intents as the new unknown OOD intents.
BANKING [55]: This intent classification dataset concerns online banking services with 77 intents. We randomly select 57 intents as the IND intents and 20 intents as the OOD intents.
SNIPS [56]: This is another intent classification dataset extracted from a voice assistant. Five out of seven intents are randomly selected to serve as the IND intents and remaining two intents are regarded as the OOD intents.
TABLE I: Original data composition of each dataset
Dataset num. of cate. Train Val Test
CLINC 150 15,000 3,000 4,500
BANKING 77 9,003 1,000 3,080
SNIPS 7 13,000 700 700
TABLE II: IND, OOD intent segmentation details
Dataset original number of categories num. of selected IND categories num. of selected OOD categories
CLINC 150 112 38
BANKING 77 57 20
SNIPS 7 5 2
TABLE III: Re-organized data composition
Dataset num. of cate. Train Val Test I Test II
CLINC 150 12,000 3,000 4,500 3,000
BANKING 77 7,463 1,000 3,080 1,540
SNIPS 7 12,300 700 700 700
TABLE IV: Data division usage in training/evaluations in this paper
IND part OOD part
Train Pre-training No usage
Val Pre-, T-4 training validation
Test I T-1 test, T-2 test T-2 test, T-3 test
Test II T-4 test
TABLE V: Examples of each dataset
Intent name Example
CLINC
translate what is the equivalent of, ’life is good’ in french
distance how many miles will it take to get to my destination
cancel_reservation please cancel the table for two at burger king
BANKING
verify_source_ of_funds Where can I check out the source of available funds in my account?
Refund_not_ showing_up I contacted the seller for a direct refund last week but he still hasn’t give me my money. What should I do?
cash_withdrawal_ not_recognised Hello, Please look into this matter urgently. As i have lost my wallet and i can see one withdrawal, now i do not want to loose more money.
SNIPS
BookRestaurant book a restaurant at sixteen o clock in sc
SearchScreeningEvent find me showtimes for animated movies in the neighbourhood
AddToPlaylist put this song in my funk playlist please
V-C Metrics
We use the following metrics for model evaluation. Basic concepts are provided here, with detailed calculations in Appendix B.
V-C1 For Task 1 and 4: F1 values
Since T-1 and T-4 are both classification problems, we use Micro and Macro F1 scores for evaluation.
V-C2 For Task 2: AUROC, AUPR, FPR90
As implied in Section III-C, the accuracy of IND/OOD detection is influenced by the comparison between $sco$ and different $\lambda$. To ensure a fair and standard evaluation, we use three threshold-independent classification metrics.
AUROC/AUPR: AUROC, or Area Under the Receiver Operating Characteristic curve, measures the area under the curve of true positive rate (TPR) against false positive rate (FPR) at varying $\lambda$ values. AUPR stands for Area Under Precision-Recall curve. Higher values in either metric indicate better OOD detection performance.
FPR90: A practical metric which shows the FPR value when TPR reaches 90%. Lower values indicate better performance.
V-C3 For Task 3: NMI, ARI, ACC
For T-3, we use three common metrics in clustering and intent discovery: Normalized Mutual Information (NMI), Adjusted Rand Index (ARI) and Clustering Accuracy (ACC). Simply speaking, NMI and ARI measure the similarity between two clusters, while ACC evaluates the clustering accuracy after finding the best true - pseudo-clustering label mapping. Higher values in these metrics indicate a better performance.
V-D Data usage and training/evaluation order
In this paper, since there are many tasks and data divisions involved in our evaluations, we explain the training and evaluation order here with Figure 3. A complete procedure in algorithm form can also be founded in Appendix D.
As shown in Figure 3, in the Pre-training stage, the model was trained using SCL with Train - IND data. During training, Val - IND/OOD data were used to determine the stopping point, specifically when the OOD detection AUROC on Val set reached its peak.
After pre-training, the basic intent classification performance of the initial model (T-1) was tested on Test I - IND data, while the OOD detection ability (T-2) was tested on Test I - IND/OOD data. The filtered OOD texts (refer to Section IV-B3 for the filtering strategy) were saved and sent to T-3 for new intent discovery. Finally, using the pseudo-labeled old IND data (labeled by initial model) and new “IND” data (labeled in T-3), the initial model was re-trained, and the classification performance was tested on Test II set (T-4).
Figure 3: Diagrams of the evaluation workflow and data division used in each task
V-E Implementation settings
We adopted BERT-base [8] and MPNet-base [57] as the base encoders for both our methods and baselines (expect for PTO-Sup that used GPT-2 small [35]). These encoders were chosen because BERT is the most commonly used pre-trained encoder, while MPNet, a BERT-like variation, enhances context learning ability for better sentence encoding [58]. For SCL setting, we set dropout rate $p\textit{=}0.1$, $n_{\text{views}}\textit{=}2$ for “view” generation, and $\tau\textit{=}0.1$ for calculating the loss. During the pre-training stage, we used the OOD detection AUROC on Val set to evaluate the training process for our SCL and FT methods, saving only the model with the highest $AUROC_{val}$. The training time reported in Table VI also corresponds to the training time of the best model. All experiments were conducted three times, and the results were averaged to obtain the final result. The experiments were run on a machine with an Ubuntu 18.04 operating system, an Nvidia A6000 GPU, an Intel i9-10980X CPU, and 128GB RAM.
VI Results and Analysis
VI-A Training time
TABLE VI: Pre-training time and validation result comparison
Dataset CLINC BANKING SNIPS
[…]
FT 18 4min10s 95.79 14 3min 93.81 15 3min15s 98.09
SCL (Ours) 7 1min10s 96.37 10 2min30s 94.97 12 2min40s 98.15
TABLE VII: T-1: intent classification result on Test I - IND set
Dataset CLINC BANKING SNIPS
Method/Metric Micro F1$\uparrow$ Macro F1$\uparrow$ Micro F1$\uparrow$ Macro F1$\uparrow$ Micro F1$\uparrow$ Macro F1$\uparrow$
FT-FullyConnect 97.23 97.22 87.32 85.69 98.83 98.74
SCL-NN (Ours) 97.23 97.23 94.17 94.16 98.83 98.75
TABLE VIII: T-2: IND/OOD detection result on Test I - IND/OOD set, the best result is in bold, the runner-up result is underlined
Dataset CLINC BANKING SNIPS
[…]
PTO-Sup 81.67 55.01 44.3 40min 79.36 56.44 55.0 72min 97.53 93.87 6.4 3min20s
Layer-CE 95.67 86.71 10.7 78min 92.89 81.67 20.3 47min 93.47 87.34 15.6 18min
FT-NDist 95.64 85.96 9.7 4min10s 92.73 80.12 18.9 3min 96.56 90.43 9.2 3min15s
SCL-NDist (Ours) 96.15 88.41 8.5 1min10s 93.44 82.06 15.7 2min30s 97.41 94.36 4.3 2min40s
Note: All results are calculated under ‘OOD example as class 1 (positive) and IND example as class 0 (negative)’ condition.
TABLE IX: T-3: New intent discovery result on Test I - OOD set, the best result is in bold, the runner-up result is underlined
Dataset CLINC BANKING SNIPS
[…]
MTP-KNNCL 92.36 79.70 83.60 6min10s 83.33 69.75 80.75 6min15s 78.04 85.50 96.26 6min40s
DPN 89.93 75.19 82.89 6min30s 82.51 69.04 76.50 6min 68.85 77.72 94.12 6min30s
FT-KMeans 86.81 67.71 76.27 4min10s 81.88 64.88 73.86 3min 66.16 70.29 91.98 3min15s
SCL-KMeans (Ours) 93.08 81.14 85.67 1min10s 84.44 70.64 81.16 2min30s 70.94 75.82 93.58 2min40s
We first show the model training time by SCL and CE fine-tuning in Table VI. It can be found that our model always reached the stop point earlier with higher Val AUROC scores, paving the way for a better performance in the following downstream tasks.
VI-B Result on task 1: user intent classification
The initial classification results on Test I - IND data are shown in Table VII. Both SCL and FT achieved very high accuracy on CLINC and SNIPS datasets. However, on BANKING, our method showed its strength, achieving significantly better F1 scores on both metrics. This is because, compared with the other two datasets, examples in BANKING have a longer input length while the inquiry expressions are also more diverse (as shown in Table V), which greatly increased classification difficulty. Despite this, our model maintained a high accuracy, demonstrating its robustness even with longer and noisier inputs.
VI-C Result on task 2: out-of-domain detection
The results on T-2 are shown in Table VIII. It can be observed that while FT provided solid results, our method performed better and achieved nearly all the best results. Layer-LOF that applied a layer-voting strategy also obtained slightly better results than FT on the first two datasets. However, the cost is a very long training time by the extensive computations required to measure information redundancy between layers. The random layer selection strategy also increased uncertainty, making it less practical.
On the other hand, PTO-Sup’s generative OOD detection method seemed work badly except on SNIPS. We speculate the reason as SNIPS exhibited more “background-shift” between IND/OOD examples. For instance, there are two intents belonging to IND and OOD parts shared the topic of “watching a movie”, although one was looking for movie information and another involved planning a cinema visit. Distinguishing this kind of OOD texts is a strong point of generative and textual perplexity-based methods but a weak point of CE fine-tune-based methods [36]. This is also evident from the performance drop of the other two fine-tuning-based baselines on SNIPS. Our method, however, still held strong across all situations, showing its superiority on OOD detection.
VI-D Result on task 3: new intent discovery
Table IX shows the results for Task 3 - new intent discovery. The feature optimization capabilities of SCL over CE fine-tuning are proven again by higher clustering scores on all datasets (We will provide the embedding space visualization evidence in following Section VI-F). As for the baselines, MTP-KNNCL also showed good clustering performance especially on SNIPS, where the other methods performed badly. We speculate that this is likely due to the existence of “background shift” discussed in T-2 result, where the unsupervised MLM target and semi-supervised KNN-ConL techniques used by MTP-KNNCL mitigated the risk of “over-training” common in supervised training settings. However, the cost is a slightly longer training time.
VI-E Result on task 4: continual learning
TABLE X: T-4: Continual learning result on Test II set
Dataset CLINC BANKING SNIPS
Method/Metric Micro F1$\uparrow$ Macro F1$\uparrow$ Micro F1$\uparrow$ Macro F1$\uparrow$ Micro F1$\uparrow$ Macro F1$\uparrow$
OVERALL
FT 92.65 93.37 79.55 81.69 94.79 94.86
SCL 97.70 97.69 91.75 91.79 99.14 99.14
ON IND
FT 97.53 97.48 86.29 84.33 98.80 98.80
SCL 98.14 98.10 93.53 93.55 99.30 99.29
Ref-Ori. FT model 98.30 98.31 87.11 85.93 99.40 99.40
Ref-Ori. SCL model 98.39 98.40 93.60 93.63 99.40 99.40
ON OOD
FT 89.53 85.65 82.14 78.04 92.18 91.52
SCL 98.48 98.46 92.51 92.30 99.75 99.75
(a) SCL (all test i data)
(b) FT (all test i data)
(c) MTP-KNNCL (all test i data)
(d) SCL (test i - ood data)
(e) FT (test i - ood data)
(f) MTP-KNNCL (test i - ood data)
Figure 4: T-SNE text feature space visualization result on BANKING test I - all data (top row) and only OOD data (bottom row)
Finally, we present the classification results of $ENC_{new}$ after the incremental learning in Table X. Overall, due to the excellent clustering performance on Task 3, this time our method further widened the accuracy gap with FT on all datasets. The concerning forgetting phenomenon also did not appear, as there was only a slight decrease on IND classification accuracy after the continual learning. This proves that our training strategy of using both pseudo-labeled IND and OOD texts is effective.
In addition to the continued high accuracy, we would like to highlight another training advantage of SCL over CE fine-tuning. In FT re-training, the last fully-connected layer had to be replaced and reinitialized to accommodate the new intents, requiring at least 4-5 epochs to obtain desirable results in our experiments. On the other hand, since SCL focuses on the feature optimization and label judgment is based on the nearest neighbor algorithm, it directly used the new data with extended labels for model training, offering straightforward scalability and efficiency.
In conclusion, our method demonstrated high standards of practicality and efficiency across all tasks, making it feasible to train and maintain a robust QA system quickly and easily.
VI-F Embedding space visualization
Finally, to evaluate the representation optimization effect, we conducted a t-SNE embedding space visualization on BANKING Test I set. The results are shown in Figure 4. The top row displays the embedding arrangement over all intents (77 categories), while the bottom row focuses only on the newly discovered (previous OOD, 20 categories) data.
From the visualization result on overall data distribution, it can be observed that both SCL and FT performed reasonably well, although FT mixed embeddings from different intents more frequently. On the other hand, MTP-KNNCL produced relatively larger clusters, with some clusters very close to each other, even starting to “stick” together. We consider this may be due to the training effect of KNN-ConL in MTP-KNNCL, which continuously pulled nearby similar embeddings together, resulting in a side effect on cluster aggregation.
The advantage of our proposed method is more obvious in the visualization on newly discovered intents. Without accessing OOD texts, our method still separated most of the OOD classes, with only two class clusters mixed. MTP-KNNCL also worked relatively well, but its dispersion effect of different intents was inferior to ours. Finally, FT produced a flatter embedding distribution with more severe confusion between intents, leading to the lowest clustering scores among the three methods.
To conclude, despite using less than half the time to the state-of-the-art method MTP-KNNCL and requiring fewer tuning steps, our method still produced a splendid underlying embedding space for clustering and other tasks, which is valuable in practical applications.
VII Conclusion
In this paper, we addressed the functionality and efficiency problems of current QA system, and proposed our solution to solve them and build a comprehensive system. We first defined four key tasks to complement the system’s functionality shortcomings, and then applied SCL as the unified representation learning method with minimal additional tuning to achieve the task targets. The experiment results demonstrated that although being simple and straightforward, the proposed method still achieved leading results in each task.
As this complete solution is first proposed and is still a prototype, there are still many situations needed to be considered for real deployment in the future. For example, there may also be some noisy inputs appearing, which are hard to be categorized into any intents. In such cases, developing a filtering algorithm before the intent discovery process might be necessary. In addition, to show efficiency, this time the downstream methods were all chosen to be as simple as possible, while more advanced methods could potentially yield even better results when combined with SCL.
Appendix A KMeans details
The KMeans learning process is shown in Algorithm 1.
Algorithm 1 KMeans clustering
1:Text embeddings $\{\bm{h}_{j^{\prime}}\}_{j^{\prime}=1}^{N_{ood}}$ of detected $\{\bm{x}_{j^{\prime}}\}_{j^{\prime}=1}^{N_{ood}}$ in T-2, number of clusters $K$
2:Label assignments $\{\hat{y}_{j^{\prime}}\}_{j^{\prime}=1}^{N_{ood}}$
3:Initialize $K$ cluster centroids $\{\bm{\mu}_{k}\}_{k=1}^{K}$ randomly.
4:repeat
5:     for each data point $\bm{h}_{j^{\prime}}$ do
6:         Assign $\bm{h}_{j^{\prime}}$ to the nearest cluster centroid:
$\small\hat{y}_{j^{\prime}}=\arg\min_{k\in\{1,2,\ldots,K\}}\|\bm{h}_{j^{\prime}}-\bm{\mu}_{k}\|^{2}$ (14)
7:     end for
8:     for each cluster $k$ do
9:         Update $\bm{\mu}_{k}$ as the mean of all points assigned to cluster $k$:
$\small\bm{\mu}_{k}=\frac{1}{|num(cate\ k)|}\sum\bm{h}_{j^{\prime}\rightarrow k}$ (15)
10:     end for
11:until the cluster assignments $\{\hat{y}_{j^{\prime}}\}$ do not change or the centroids $\{\bm{\mu}_{k}\}$ converge
12:return Label assignments $\{\hat{y}_{j^{\prime}}\}_{j^{\prime}=1}^{N_{ood}}$
Appendix B Metric calculation details
Here we show the detailed calculation process of the used metrics.
B-A For Task 1 and 4: F1 values
The Micro and Macro F1 scores are calculated as follows:
Micro F1 score: is calculated by summarizing the classification situations over all classes. It is the harmonic mean of micro-averaged precision ($P$) and recall ($R$):
[…]
Macro F1 score: calculates the F1 score independently for each class and then takes the average, which treats all classes equally:
[…]
where $C$ is the number of categories.
B-B For Task 3: NMI, ARI, ACC
The NMI, ARI, ACC are calculated as follows:
Normalized Mutual Information (NMI): is a measure of the similarity between two sets of clusters:
$\text{NMI}(U,V)=\frac{2\cdot I(U;V)}{H(U)+H(V)}$ (18)
where $I(U;V)$ is the mutual information between clusters $U$ and $V$, and $H(U)$ and $H(V)$ are the entropies of $U$ and $V$ respectively. It ranges from 0 (no mutual information) to 1 (perfect correlation).
Adjusted Rand Index (ARI): measures the similarity between two data clusterings by considering all pairs of samples, and counting pairs that are assigned in the same or different clusters in the predicted and true clusterings:
$\text{ARI}=\frac{\sum_{ij}\binom{n_{ij}}{2}-\left[\sum_{i}\binom{a_{i}}{2}\sum_{j}\binom{b_{j}}{2}\right]/\binom{n}{2}}{\frac{1}{2}\left[\sum_{i}\binom{a_{i}}{2}+\sum_{j}\binom{b_{j}}{2}\right]-\left[\sum_{i}\binom{a_{i}}{2}\sum_{j}\binom{b_{j}}{2}\right]/\binom{n}{2}}$ (19)
where $n_{ij}$ is the number of elements in the intersection of clusters $i$ and $j$, $a_{i}$ is the sum over row $i$, and $b_{j}$ is the sum over column $j$.
Clustering Accuracy (ACC): evaluates the clustering result after finding the best one-to-one mapping between cluster labels and true labels:
$\text{ACC}=\frac{\sum_{i=1}^{n}\delta(y_{i},\text{map}(\hat{y}_{i}))}{n}$ (20)
where $y_{i}$ is the true label, $\hat{y}_{i}$ is the cluster label, $\text{map}()$ is the mapping function where we use the Hungary algorithm this time, and $\delta$ is the Kronecker delta function.
Appendix C Abbreviations
Table XI shows the abbreviations that have appeared in this paper and their complete names.
TABLE XI: Abbreviations and full names
Abbreviations Full name
QA question answering (system)
RepL representation learning (feature learning)
ConL contrastive learning
MeL metric learning
UConL unsupervised contrastive learning
SCL supervised contrastive learning
IND in-domain (texts)
OOD out-of-domain (texts)
GDA Gaussian discriminant analysis
LOF local outlier factor
GPT Generative Pre-Trained Transformer
BERT Bidirectional Encoder Representations from Transformers
CE cross-entropy (loss)
FT fine-tuning
MLM masked language modeling
NN nearest neighbor
Appendix D Learning and evaluation procedure (Algorithm form)
The learning and evaluation procedures of pre-training and each task are shown in Algorithm 2.
Algorithm 2 Overall process
1:Train data, Val data, Test I data, Test II data
2:Solutions for T-1 to T-4
3:      -Pre-training Stage-
4:$Epoch\leftarrow 0$, $AUROC_{best-val}\leftarrow 0$
5:while $Epoch<20$ do
6:     Train ENC on $\bm{x}_{\text{train}}$ using SCL loss object (eq. 8)
7:     Calculate OOD detection $AUROC_{val}$ on this epoch
8:     if $AUROC_{val}>AUROC_{best-val}$ then
9:         Save model of this epoch, and record the threshold $\lambda$ when TPR=90%
10:     end if
11:end while
12:        -Task 1-
13:Input $\bm{x}_{\text{test I}}$ (IND part) to ENC to obtain the representations:
14:    $\bm{h}_{\text{test I}}=ENC(\bm{x}_{\text{test-1}})$
15:The label is determined as the label of the nearest IND cluster centroid:
16:    $\hat{y}_{\text{test I}}=\arg\min_{c\in\mathcal{C}}MDist(\bm{h}_{\text{test I}},\bar{\bm{h}}_{c})$
17:Calculate Micro, Macro F1 values
18:        -Task 2-
19:Input $\bm{x}_{\text{test-1}}$ (IND, OOD part) to ENC to obtain the representations and calculate the detection scores:
20:    $sco(\bm{x}_{\text{test I}})={\underset{c\in{\mathcal{C}}}{min}}\ MDist(\bm{h}_{\text{test I}},\bar{\bm{h}}_{c})$
21:Calculate FPR, TPR results on various thresholds $\lambda$, and finally AUROC, AUPR, FPR90 values
22:      -Task 3 preparation-
23:Determine $\lambda_{filter}$ when TPR=90% on $\bm{x}_{\text{val}}$
24:Divide $\bm{x}_{\text{test-1}}$ into $\bm{x}_{\text{test-1, IND}}$ and $\bm{x}_{\text{test-1, OOD}}$ two parts by comparing detection $sco$ and $\lambda_{filter}$
25:        -Task 3-
26:Apply KMeans on $\bm{x}_{\text{test-1, OOD}}$ and give new pseudo labels to $\bm{x}_{\text{test-1, OOD}}$ as $\hat{y}_{\text{test-1, OOD}}\in\mathcal{C}_{kn^{\prime}}$
27:Calculate NMI, ARI, ACC metrics
28:        -Task 4-
29:Give pseudo labels to $\bm{x}_{\text{test-1, IND}}$ using the initial ENC as $\hat{y}_{\text{test-1, IND}}\in\mathcal{C}_{kn}$
[…]
31:Test on $\bm{x}_{\text{test II}}$ and calculate Micro, Macro F1 values
32:END
[…]