### v2

Abstract
This paper presents a novel and comprehensive solution to enhance both the robustness and efficiency of question answering (QA) systems through supervised contrastive learning (SCL). Training a high-performance QA system has become straightforward with pre-trained language models, requiring only a small amount of data and simple fine-tuning. However, despite recent advances, existing QA systems still exhibit significant deficiencies in functionality and training efficiency. We address the functionality issue by defining four key tasks: user input intent classification, out-of-domain input detection, new intent discovery, and continual learning. We then leverage a unified SCL-based representation learning method to efficiently build an intra-class compact and inter-class scattered feature space, facilitating both known intent classification and unknown intent detection and discovery. Consequently, with minimal additional tuning on downstream tasks, our approach significantly improves model efficiency and achieves new state-of-the-art performance across all tasks.
Index Terms:
Intent discovery, natural language processing, out-of-domain detection, question-answering system, supervised contrastive learning, text classification.
[…]
=-21pt
I Introduction
Figure 1: The workflow of the proposed comprehensive and adaptive QA system. First, an initial model is trained using current known data (Pre). Then, during daily usage, an inquiry will first be judged to be an answerable input or not (T-2): if so, proceed to the intent detection (T-1) and retrieve the answer; if not, the unknown inquiry will be temporally saved and be clustered later (T-3). Finally, the initial model is re-trained with newly discovered data to obtain the classification ability to new intents (T-4).
Question answering (QA) system is a well-established and important application widely used in areas such as healthcare [1], finance [2] and e-commerce [3]. Despite the world gradually moving into the era of large language models (LLMs), classic QA systems still maintain a strong presence due to their low cost, ease of deployment, and high accuracy in answering questions.
A typical QA system mainly comprises two components: user intent classification and answer retrieval. The efficiency of the former is crucial, as a correct response often depends on accurately determining the user’s intent. To achieve robust classification, it is essential to efficiently extract features from user input, providing intent-specific and discriminative text representations for classification. Initially, text features were extracted manually using techniques like term frequency-inverse document frequency (TF-IDF [4]) or bag-of-words (BoW [5]). With the advent of neural networks, feature extraction became an automated process [6, 7]. Today, the power of pre-trained Transformers like BERT [8] has made it much easier to extract features and build classification models, requiring only a small amount of data with Cross-Entropy (CE) loss fine-tuning [9, 10].
It seems that the current QA system building is simple and mature, with nearly nothing to be changed. However, from the perspective of building a comprehensive system, current QA systems still face unresolved issues at both internal and external levels, which have not been fully addressed yet:
  1. $\bullet$
    Externally, or at the system level: Current QA systems typically adopt a “training-as-complete” policy, meaning that once the model has finished training, its knowledge is not updated. This approach leaves the system vulnerable to unseen, unknown inputs (as out-of-domain, OOD inputs [11]). To maintain the robustness and security of the model, there should be mechanisms to handle these inputs. Furthermore, even when new data are collected to upgrade the system, the common practice is to re-train a new model from scratch, leading to a waste of labor and resources.
  2. $\bullet$
    Internally, or in training efficiency: Increasing studies indicate that the BERT+CE training combination still has efficiency problems. For BERT, the raw text feature distribution is anisotropic [12]. Additionally, when using CE as the fine-tuning loss, text features of different intents often remain mixed with unclear boundaries in the embedding space [13]. Both issues hinder training efficiency and model performance.
In this paper, aiming to build a more robust QA system, we propose a novel and comprehensive solution to address both issues. To tackle the systematic issue, including the basic intent classification function, we define four key tasks and design a workflow between them (as shown in Figure 1), enabling the system to be self-consistent and recurrently updatable. Specifically, the tasks are:
  1. T-1
    User input intent classification: The fundamental task for a QA system.
  2. T-2
    OOD input detection: Detect abnormal inputs to prevent misleading output. The detected OOD texts are saved as source data for Task 3.
  3. T-3
    New intent discovery: Utilize detected OOD texts to discover new possible intents once they accumulate to a sufficient quantity.
  4. T-4
    Continual learning: Use newly discovered and pseudo-labeled data to incrementally train the initial classification model, enhancing the system’s capability to handle the latest inquiries.
To address the training efficiency problem, we leverage supervised contrastive learning (SCL, [14]) to replace CE fine-tuning as the unified text representation optimization method, which requires minimal tuning to achieve all four task objectives. We choose SCL because, in our previous research on OOD detection [15], it demonstrated excellent capability in upstream feature optimization, allowing an accurate OOD text detection using only a simple distance-based algorithm in downstream.
Figure 2: Diagrams of SCL’s representation optimization procedure, and solution for each task. Left: the training and optimization effect by SCL, which gathers same-intent text features together and pushes different features further apart, providing an underlying embedding space for following tasks. Right: the classification/detection processes of T-1 to T-4 after SCL. (Together with Figure 1 for better understanding)
More specifically, the advantage of SCL over CE fine-tuning lies in its ability to enhance both learning efficiency and feature separation. As shown in Figure 2-b, SCL employs a simple but clear optimization target: pulling same-label text features together and pushing different ones further apart, resulting in an intra-class compact and inter-class scattered, well-organized embedding space. This clear separation between clusters facilitates easier in-domain (IND) intent classification and IND/OOD separation and detection (Figure 2, T-1/2). Additionally, this structure also supports the aggregation of currently unknown but semantically similar features, creating a conducive environment for new intent discovery and continual learning (Figure 2, T-3/4). Consequently, even with the simplest algorithms in downstream tuning, all task objectives can be achieved effortlessly and efficiently.
Our contributions can be summarized as follows:
  1. 1.
    For the first time, we clearly identify the systematic and efficiency issues in current QA systems and propose a comprehensive solution involving four key tasks in a natural workflow to enhance the training process, resulting in a robust and comprehensive system.
  2. 2.
    To achieve the task objectives, we leverage SCL as the unified representation learning approach, which requires minimal tuning in downstream tasks.
  3. 3.
    Our experiments demonstrate that, despite its conciseness and simplicity, the proposed solution significantly improves global feature optimization efficiency and outperforms state-of-the-art methods across all tasks, complementing the deficiencies of current QA systems. [^1] Code is available at https://github.com/bowang-rw-02/scl-one-for-four
II Related Works
II-A Representation learning and intent classification model building in QA system
As introduced in Section I, the efficiency of a QA system heavily depends on the performance of the intent classification module, where the representation learning (RepL) process plays a key role in it. Here, a good RepL method can capture the intrinsic patterns and structures of the data, and thus providing a solid foundation for downstream classification and other tasks. Early RepL research focused on manually extracting meaningful features from raw data, with preliminary methods including TF-IDF [4] and BoW [5]. Subsequently, the development of word embeddings extracted via neural networks automated the feature extraction process [6, 7]. More recently, with the advent of large-scale language encoders like BERT [8] that are pre-trained on extensive corpora, obtaining high-quality text embeddings has become much easier. The common classification model building approach is to use BERT as the base encoder, with CE fine-tuning as the RepL method, which is widely used in various recent QA systems [16, 10, 1, 17].
Drawback of BERT+CE mode
However, this BERT+CE paradigm is proven to be suboptimal for feature optimization these years. Firstly, from a global perspective, [18, 12] reported that the overall text embedding distribution generated by BERT is anisotropic. This means that most embeddings are spatially distributed in an elongated cone-like space rather than uniformly in a sphere, leading to high similarity between any two sentence representations and thus posing challenges for downstream tasks. Secondly, from a local perspective, CE fine-tuning usually results in a “loose” feature arrangement [13, 14], where the boundaries between different intent clusters are barely distinguishable. This not only reduces the performance of known, IND intent classification, but also makes the model prone to confusing IND inputs with future unseen but similar OOD inputs.
Contrastive learning as a RepL method
To overcome above limitations and further improve learning efficiency, the contrastive learning (ConL), has been actively studied recently. Earliest ConL can be traced back to supervised metric learning (MeL) [19] used in image classification tasks. Different with the vague and loose optimization goal of CE fine-tuning mode, MeL applies a simpler but straightforward target of “gathering similar examples together and pushing dissimilar ones further apart”. This helps generate compact feature clusters efficiently, and established the foundation of ConL. Subsequently, unsupervised ConL (UConL) [20, 21] was proposed, sharing the same strategy but aiming to improve the overall representation uniformity. Finally, SCL [14] incorporates the advantages of both MeL and UConL, thus having the potential to address both global and local feature optimization issues.
Despite these advantages, currently ConL is still mainly examined in computer vision (CV) field. In natural language processing (NLP) field, there are only limited applications, primarily in natural language inference (NLI) tasks using UConL methods [22, 23]. A boarder exploration of SCL’s learning effect across various tasks, including its application in QA systems, remains unaddressed, which is exactly our main research target this time.
II-B Past work on individual tasks
To provide a boarder perspective, we briefly review previous methods for addressing the three newly introduced tasks.
T-2 Out-of-Domain, OOD detection: Although QA systems have evolved significantly in recent years, they can only work correctly on the known inquires and lack the ability to recognize and reject unknown inputs, which raises the issue of OOD detection [11]. To solve this problem, early methods built a binary IND/OOD detection system using numerous OOD data [24]. However, the scope and variety of OOD data can be infinite, making this method impractical. Subsequently, more approaches have focused on training networks using only IND data to generate heterogeneous representations [25, 26], then applied various downstream algorithms, such as local outlier factor (LOF) [27], Gaussian discriminant analysis (GDA) [28], energy score [29, 30], and Mahalanobis distance [31, 32], to distinguish IND and OOD text features. Latest advances fall into two broad categories: using ConL methods to optimize the representation space [33, 34, 15], or employing generative modeling approaches like training a GPT model [35] on IND data to achieve lower likelihoods for OOD text generation [36, 37, 38].
T-3 New intent discovery, NID: NID is a classic and substantially studied task in NLP. Early studies focused on identifying new categories directly from unlabeled data, with most methods applying unsupervised clustering [39, 40]. The later approaches commonly incorporated some labeled data to support the discovery of unknown intents [41, 42]. A typical workflow involves computing textual similarities between unlabeled and known data, assigning pseudo labels to unknown data based on the results, and then performing clustering. Recently, NID research has also emphasized the importance of RepL, with many studies incorporating pre-training and contrastive learning methods to optimize feature space arrangements [43, 44, 45].
T-4 Continual learning: Continual learning generally denotes a re-training procedure to let the network learn new knowledge in stages, and finally reach a level to handle all tasks. For example, we may let a model learn the even five handwriting numbers and then the odd fives, finally let it recognize all ten numbers. The biggest challenge of continual learning is catastrophic forgetting [46], where the model’s performance on earlier tasks deteriorates as it learns new tasks. Current studies have also been focusing on solving this issue, where the solutions mainly include: a) weight regularization [47, 48], which imposes some constraints on the model’s parameters to preserve values of model weights associated with knowledge from old data; b) knowledge distillation [49, 50], which builds a smaller model to mimic the past model output as the hint for old tasks during handling new tasks; c) replay [51, 52], which involves storing a subset of old data and replaying it during training on new tasks to reinforce the old knowledge.
Although above methods have made much progress in their respective tasks, there still remains a gap in systematically integrating these tasks into a comprehensive QA system building, which is the aim of this paper.
III Task and workflow definition
As mentioned in Section I, we introduce four basic tasks to build a comprehensive and adaptive QA system (Figure 1). Before presenting the specific solutions in the next section, we formally define these tasks and also simply introduce the workflow for better comprehension.
III-A Preliminary
[…]
Encoder: We have a language encoder ENC that encodes and transforms a discrete text $\bm{x}$ into an m-dimension continuous representation $\bm{h}$: $\bm{x}$$\mathbb{R}^{m}$. Since there are many text similarity calculations during SCL training, we adopt a simple but effective “mean-pooling” strategy to obtain the sentence representation. Specifically, the average of all word embeddings in the sentence is served as the sentence embedding:
$\small\bm{h}=\frac{1}{len(\bm{x})}\sum_{i=1}^{len(\bm{x})}ENC(w_{i})$ (1)
where $len(\bm{x})$ is the length of input $\bm{x}$, $\{w_{1},...,w_{i},...,w_{len(\bm{x})}\}$ are the words in the sentence $\bm{x}$.
III-B T-1: User intent classification
The target of this task is to find an accurate mapping:
$f_{cls}:ENC(\bm{x}_{i})\mapsto\hat{y}_{i}\qquad\hat{y}_{i}\in\mathcal{C}_{kn}$ (2)
For traditional methods like CE-fine-tuning, this is achieved by further adding a fully-connected layer to the end of ENC and training them together using CE loss. In the proposed method, ENC is pre-trained using SCL (refer to Section IV-A for details) and the parameters are fixed during T-1 and the following T-2 and T-3. The $f_{cls}$ is realized by a nearest neighbor algorithm (Section IV-B1).
III-C T-2: Out-of-Domain detection
The target of this task is to determine whether an example $\bm{x}_{j}$ from $\mathcal{S}_{unlabeled}$ belongs to $\mathbb{D}_{IND}$ or $\mathbb{D}_{OOD}$. This is usually realized by comparing the detection score $sco$ of $\bm{x}_{j}$ with an empirically-set threshold $\lambda$, where $sco$ is given by a detection algorithm $f_{det}$ analyzed on the text embedding:
$sco_{j}=f_{det}(ENC(\bm{x}_{j}))$ (3)
$\bm{x}_{j}\in\left\{\begin{array}[]{l}\mathbb{D}_{OOD},\;if\;sco_{j}\geq\lambda\\ \mathbb{D}_{IND},\;if\;sco_{j}<\lambda\end{array}\right.$ (4)
In the actual workflow (as shown in Figure 1), the detected $\bm{x}_{j^{\prime}}^{ood}$ samples are saved for T-3 (new intent discovery), while the detected $\bm{x}_{j^{\prime}}^{ind}$ examples will be classified normally in T-1 and also saved for the final task, T-4 (continual learning).
III-D T-3: New intent discovery
Once the collected OOD texts $\bm{x}_{j^{\prime}}^{ood}$ from T-2 reach a sufficient quantity, an algorithm $f_{nid}$ can be used to cluster and assign pseudo labels $\hat{y}_{j^{\prime}}$ to these examples, as the “new IND intent” labels stored in set $\mathcal{C}_{kn^{\prime}}$:
$f_{nid}:ENC(\bm{x}_{j^{\prime}}^{ood})\mapsto\hat{y}_{j^{\prime}}\qquad\hat{y}_{j^{\prime}}\in\mathcal{C}_{kn^{\prime}}$ (5)
III-E T-4: Continual learning
[…]
$f_{cls\text{-}new}:ENC_{new}(\bm{x}_{k})\mapsto\hat{y}_{k}\qquad\hat{y}_{k}\in\{\mathcal{C}_{kn}+\mathcal{C}_{kn^{\prime}}\}$ (6)
IV Methodology
In this section, we introduce the methodology of SCL as the RepL method in the pre-training stage, followed by the downstream methods to solve to each task.
IV-A Stage 1: Supervised contrastive learning -based pre-training
IV-A1 Principle and optimizing effect of SCL
As discussed in Section II, BERT+CE fine-tuning exhibits issues with local loose borders and a global anisotropic distribution in RepL, while SCL has the potential to solve both of these problems.
Specifically, as illustrated in Figure 2-b, for the local aspect, compared with CE fine-tuning, SCL applies a more straightforward yet effective target: pulling the representations of similar examples (as positive examples) together and pushing dissimilar ones (as negative examples) further apart. To tackle the global distribution problem, SCL further generates “view” augmentations as additional positive examples before the contrasting process (Figure 2-Pre-1), optimizing features in a boarder scale. Together, these strategies provide a better supervision signal to arrange text representations to form compact and distinct clusters, enhancing both the alignment and uniformity of the feature space, and thus sharpening the distinctions between texts from various intents.
This well-organized embedding space provides a strong foundation that benefits all downstream tasks (as Figure 2-right part shows):
  1. T-1:
    The clear separations between each intent cluster facilitate easy and accurate classification.
  2. T-2:
    The compact and distinct IND clusters highlight the differences between them and OOD texts that do not belong to any known intents.
  3. T-3:
    Although not trained directly with OOD data, the two optimization directions of SCL still drive the aggregation of semantically similar features. This creates an environment conducive to intent discovery algorithms, making it easier to identify the pre-gathered data.
  4. T-4:
    The accurately discovered new intent data, pre-organized feature space, and SCL-based re-training, together ensure the continuous high-performance of the updated system.
IV-A2 Generation of text “views”
As discussed in the last subsection, the effective optimization of SCL is attributed to both the supervised positive-negative movement and the unsupervised “view” augmentation. This augmentation also serves as a precaution against the lack of positive examples in the mini-batch, especially when the data quantity is small compared with category numbers.
For images, view generation can be achieved by simple transformations such as rotating and cropping. However, these techniques are not easily applicable to text “view” generation, as hasty word swaps or deletions can distort the original meaning. In our previous study [15], we found that encoding a sentence twice with different drop-out masks (Figure 2-Pre-1) is an effective way to generate high-quality, slightly different but main meaning preserved text embeddings. By using this generation method, SCL gave the best OOD detection results compared to no-view-generation [33] or more complex techniques like embedding-attack [53] (refer to Tables 3-5 in [15] for detailed results). Considering these advantages, we continue to use this technique as the augmentation method in this paper.
IV-A3 Formal learning procedure definition
Here we give the formal definition of an SCL procedure:
Step 1: Text “View” generation
[…]
$\bm{h}_{2i^{\prime}-1}=ENC_{drop1}(\bm{x}_{i}),\quad\bm{h}_{2i^{\prime}}=ENC_{drop2}(\bm{x}_{i})$ (7)
[…]
Step 2: Supervised contrastive learning
The SCL training loss is defined as follows:
$\small{\mathcal{L}}_{SCL}\text{=}\sum_{i^{\prime}\text{=}1}^{2N}\frac{-1}{\left|Pos(i^{\prime})\right|}\sum_{p\in Pos({i^{\prime}})}\log\frac{exp(\bm{h}_{i^{\prime}}\cdot\bm{h}_{p}/\tau)}{{\displaystyle\sum_{q\in Neg({i^{\prime}})}}exp(\bm{h}_{i}^{\prime}\cdot\bm{h}_{q}/\tau)}$ (8)
[…]
IV-B Stage 2: Solution for each task after SCL
With a well-organized embedding space achieved through SCL, now all four downstream tasks can be solved using simple methods with the least efforts.
IV-B1 Solution for T-1 intent classification
For T-1, as shown in Figure 2-T-1, we apply a Mahalanobis distance (MDist) - nearest neighbor (NN)-based method to classify the text representations encoded by SCL-trained ENC, which does not require any additional training. We choose MDist because, compared to Euclidean distance, MDist considers the distribution of each feature component, providing a more accurate measurement of text similarity. Additionally, to further accelerate the classification speed but avoid the instability of NN, we propose calculating MDist between the current test sample $\bm{x}_{test}$ and only the centroids $\bar{\bm{x}}_{c}$ of each intent $c$. Finally, the label of one test example $y_{test}$ is determined as the label of the nearest centroid. The whole process can be described into the following three steps:
a) Class centroid calculation
The class centroid of category c is the average data point of all text embeddings belonging to this intent:
$\bm{h}_{c,i}=ENC(\bm{x}_{c,i})$ (9)
$\small\bar{\bm{h}}_{c}=\frac{1}{num(c)}\sum_{i=1}^{num(c)}\bm{h}_{c,i}$ (10)
where $num(c)$ is data quantity of category c, and $\bm{h}_{c,i}$ is the text representation of its i-th example $\bm{x}_{c,i}$.
b) MDist calculation
We calculate the covariance matrix $S$ over all data (eq. 11) and then calculate MDist (eq. 12).
$\small S=\frac{1}{\left|{\mathcal{C}}_{kn}\right|}\sum_{c\in{\mathcal{C}_{kn}}}(\bm{H}_{c}-\overline{\bm{H}}_{c}){(\bm{H}_{c}-\overline{\bm{H}}_{c})}^{T}$ (11)
$MDist(\bm{h}_{test},\bar{\bm{h}}_{c})=\sqrt{{(\bm{h}_{test}-\bar{\bm{h}}_{c})}^{T}S^{-1}(\bm{h}_{test}-\bar{\bm{h}}_{c})}$ (12)
where $\bm{H}_{c}\text{=}[\bm{h}_{c,1},\bm{h}_{c,2},...,\bm{h}_{c,num(c)}]^{T}$ is the sample matrix of category $c$, $\overline{\bm{H}}_{c}\text{=}[1,1,..,1]^{T}\bar{\bm{h}}_{c}$.
c) Label determination
Finally, the label is determined by the nearest centroid:
$\small y_{\text{test}}=\arg\min_{c\in\mathcal{C}_{kn}}MDist(\bm{h}_{\text{test}},\bar{\bm{h}}_{c})$ (13)
IV-B2 Solution for T-2 out-of-domain detection
The method to detect OOD examples is also very simple: we still calculate the embedding distance MDist between sample $\bm{x}_{test}$ and the nearest intent centroid as T-1, but the distance is used directly as the detection $sco$:
$sco(\bm{x}_{test})={\underset{c\in{\mathcal{C}}_{kn}}{min}}\ MDist(\bm{h}_{\text{test}},\bar{\bm{h}}_{c})$ (14)
This text representation distance itself can serve as the $sco$ and be an effective detection indicator, since a larger value means that even taking the shortest distance, current example is still far from, or say has less similarity to any known, IND category centroids. As shown in Figure 2-T-2, this suggests that the current $\bm{x}_{test}$ is more likely to be an unseen, i.e., OOD text. In practical usage, as introduced in task definition (Section III-C), this $sco$ is further compared with a threshold $\lambda$, and texts with an exceeding distance will be judged as OOD examples.
IV-B3 Solution for T-3 new intent discovery
To conduct new intent discovery, the first step is to obtain the unknown data to be clustered from T-2. We achieve this by first finding a threshold $\lambda$ with a policy that when the model in T-2 gets a true positive rate, TPR=90% on validation data (i.e., 90% OOD examples in validation set are correctly recognized), then we filter out OOD texts by comparing $sco$ with this $\lambda$.
Next, since SCL is able to provide a well-structured feature space, we do not conduct any further feature pre-processing or analysis but directly apply KMeans, one of the simplest unsupervised clustering method, as T-3’s solution. Taking the representations of detected OOD samples $\bm{x}_{j^{\prime}}^{ood}$, KMeans clusters and assigns pseudo-labels $\hat{y}_{j^{\prime}}$ to them. The detailed clustering procedure of KMeans in algorithm is provided in Appendix A. Finally, the pseudo-labeled new intent examples $\{\bm{x}_{j^{\prime}}^{ood},\hat{y}_{j^{\prime}}\}$ will serve as the training data for the final task T-4.
IV-B4 Solution for T-4 continual learning
[…]
This training method can be regarded as a straightforward “replay” method to mitigate forgetting of the old intents. We choose this strategy because the goal of continual learning in our QA system is to enable the model to classify all categories (both old and new intents) simultaneously, rather than in separate stages like other studies. Additionally, since the old intent data are readily available and can be stored without significant cost, as a result, the replay strategy is simple, easy-to-deploy and well-suited to our objectives. Experiment results (Section VI-D) also demonstrate that this replay policy effectively prevents forgetting, with only minimal performance degradation observed in classifying old data.
V Experiment
V-A Datasets, division, and evaluation order
To clearly present the evaluation process for the four tasks, we begin by introducing the original dataset, then specify the data divisions applied for the experiment, and finally outline the evaluation workflow based on these divisions.
V-A1 Dataset
We use three popular public intent classification datasets for the experiment.
CLINC [54]: This is a widely-used dataset in intent classification and OOD detection tasks. It provides 150 intents of texts and one OOD intent data (that is, on OOD detection task, the former 150 intents can be treated as the IND intents). Because its OOD part does not contain specific labels that are necessary for T-3 and T-4 evaluation, we only use its 150-intent part data this time.
BANKING [55]: This is an intent classification dataset concerning online banking services, which contains around 13,000 records in 77 intents.
SNIPS [56]: This is another intent classification dataset extracted from a voice assistant, consisting of around 14,000 records in 7 intents.
The original data composition of above datasets is shown in Table I. Some examples are also provided in Table II. Since all data are well-formed, which keep a relatively good data balance and do not contain personal information, we did not take further pre-processing steps but used them as they are.
TABLE I: Original data composition of each dataset
Dataset num. of cate. Train Val Test
CLINC 150 15,000 3,000 4,500
BANKING 77 9,003 1,000 3,080
SNIPS 7 13,000 700 700
TABLE II: Examples of each dataset
Intent name Example
CLINC
translate what is the equivalent of, ’life is good’ in french
distance how many miles will it take to get to my destination
cancel_reservation please cancel the table for two at burger king
BANKING
verify_source_ of_funds Where can I check out the source of available funds in my account?
Refund_not_ showing_up I contacted the seller for a direct refund last week but he still hasn’t give me my money. What should I do?
cash_withdrawal_ not_recognised Hello, Please look into this matter urgently. As i have lost my wallet and i can see one withdrawal, now i do not want to loose more money.
SNIPS
BookRestaurant book a restaurant at sixteen o clock in sc
SearchScreeningEvent find me showtimes for animated movies in the neighbourhood
AddToPlaylist put this song in my funk playlist please
V-A2 Data division for the experiment
We apply two key divisions to adapt the data for the experimental requirements. The first is the IND/OOD intent split, where approximately 75% number of the original intents are randomly selected as the known IND intents, with the remaining 25% treated as the unknown OOD intents. Specifically, in CLINC, BANKING and SNIPS datasets, 112, 57, and 5 intents were randomly selected as the known IND intents, respectively, while the remaining 38, 20, and 2 intents were designated as the unknown OOD intents.
Additionally, since T-4 requires the recognition results from T-2 and T-3 for the final re-training (refer to Section III-E for this process and Figure 3 for workflow), the original single test set is insufficient to evaluate all four tasks, especially as it may cause data leakage for T-4. To address this, we further split around 20% data from the original training set to create a secondary test set, Test II. The new composition of the re-divided data is presented in Table III.
TABLE III: Re-organized data composition
Dataset IND/OOD cate. Train Val Test I Test II
CLINC 112/38 12,000 3,000 4,500 3,000
BANKING 57/20 7,463 1,000 3,080 1,540
SNIPS 5/2 12,300 700 700 700
V-A3 Evaluation order and data usage
Here we introduce the training and evaluation processes using above divided data. A simple workflow diagram is shown in Figure 3, while the complete procedure in pseudo-code form can also be found in Appendix D.
Training stage: As shown in Figure 3, in the pre-training stage, the model is trained with an SCL target using only the IND part of training data (Train-IND). During training, Val - IND/OOD data are used to determine the stopping point, specifically when the OOD detection AUROC (refer to Section V-B for definition) on Val set reached its peak.
Evaluation stage: After pre-training, the intent classification performance (T-1) of the initial model is first tested on Test I-IND data. The OOD detection ability (T-2) is tested on Test I-IND/OOD data. After that, the filtered OOD texts from T-2 (refer to Section IV-B3 for filtering strategy) are used for new intent discovery (T-3) and the accuracy is then evaluated. Finally, using the pseudo-labeled old IND data (labeled by initial model) and new “IND” data (labeled in T-3), the initial model is re-trained (T-4), and the classification performance is tested on Test II set.
Figure 3: Diagram of the training and evaluation workflow
V-B Metrics
We use the following metrics for model evaluation in each task. Basic concepts are provided here, with detailed calculation processes in Appendix B.
V-B1 For Task 1 and 4: F1 values
Since T-1 and T-4 both belong to classification problems, we use Micro and Macro F1 scores for the evaluation.
V-B2 For Task 2: AUROC, AUPR, FPR90
As implied in Section III-C, the accuracy of IND/OOD detection is influenced by the comparison between $sco$ and different $\lambda$. To ensure a fair and standard evaluation, we use three threshold-independent classification metrics.
AUROC/AUPR: AUROC, or Area Under the Receiver Operating Characteristic curve, measures the area under the curve of true positive rate (TPR) against false positive rate (FPR) at varying $\lambda$ values. AUPR stands for Area Under Precision-Recall curve. Higher values in either metric indicate better OOD detection performance.
FPR90: A practical metric which shows the FPR value when TPR reaches 90%. Lower values indicate better performance.
V-B3 For Task 3: NMI, ARI, ACC
For T-3, we use three common metrics in clustering and intent discovery: Normalized Mutual Information (NMI), Adjusted Rand Index (ARI) and Clustering Accuracy (ACC). Simply speaking, NMI and ARI measure the similarity between two clusters, while ACC evaluates the clustering accuracy after finding the best true - pseudo-clustering label mapping. Higher values in these metrics indicate a better performance.
V-C Baselines
We apply the following latest and competitive methods as baselines for evaluation.
Fine-tuning, FT (For all four tasks): Despite being the most common training method, CE-based fine-tuning remains one of the most powerful methods actively used among all tasks [16, 31, 41]. Thus this time we applied it as a baseline of RepL, and tested the effects throughout all four tasks. After fine-tuning, the downstream detection and clustering methods are all the same as SCL.
PTO-Sup [38] (For T-2): This work proposed a Prefix Tuning-based OOD detection method, training a generative GPT-2 model on IND data and then using the likelihood as $sco$. It assumes that the unseen OOD text should let model output a lower likelihood. This method originally supports unsupervised learning, but for a fair comparison, we applied its supervised mode with IND text labels available.
Layer-LOF [27] (For T-2): This paper addressed that pre-trained language models like BERT may over-rationalize the input. That is, even it is an OOD input, the last-layer embedding may tend to be more like IND. Based on this observation, it proposed to consider a decision of multiple layers after fine-tuning, and finally used local outlier factor (LOF) with a voting mechanism for the final decision. In this paper, we simply averaged its best-four layer results as the final result.
MTP-KNNCL [44] (For T-3): The state-of-the-art NID method that first pre-trains the model with a Masked Language Modeling (MLM) target to obtain prior semantic knowledge, then performs KNN contrastive learning for better cluster aggregation. The downstream clustering method is KMeans.
DPN [45] (For T-3): This Decoupled Prototypical Network first initializes some prototypes for known and unknown intents and aligns them as a bipartite matching problem. Unaligned prototypes represent new intents, and all embeddings are further optimized using semantic-aware prototypical learning. The downstream method is also KMeans.
TABLE IV: T-1: intent classification result on Test I - IND set
Dataset CLINC BANKING SNIPS
Method/Metric Micro F1$\uparrow$ Macro F1$\uparrow$ Time/Epoch Micro F1$\uparrow$ Macro F1$\uparrow$ Time/Epoch Micro F1$\uparrow$ Macro F1$\uparrow$ Time/Epoch
FT-FullyConnect 97.23 97.22 4min10s/18 87.32 85.69 3min/14 98.83 98.74 3min15s/15
SCL-NN (Ours) 97.23 97.23 1min10s/7 94.17 94.16 2min30s/10 98.83 98.75 2min40s/12
TABLE V: T-2: IND/OOD detection result on Test I - IND/OOD set, the best result is in bold, the runner-up result is underlined
Dataset CLINC BANKING SNIPS
[…]
PTO-Sup 81.67 55.01 44.3 40min/8 79.36 56.44 55.0 72min/16 97.53 93.87 6.4 3min20s/13
Layer-LOF 95.67 86.71 10.7 78min/19 92.89 81.67 20.3 47min/12 93.47 87.34 15.6 28min/8
FT-NDist 95.64 85.96 9.7 4min10s/18 92.73 80.12 18.9 3min/14 96.56 90.43 9.2 3min15s/15
SCL-NDist (Ours) 96.15 88.41 8.5 1min10s/7 93.44 82.06 15.7 2min30s/10 97.41 94.36 4.3 2min40s/12
Note: All results are calculated under ‘OOD example as class 1 (positive) and IND example as class 0 (negative)’ condition to highlight the model’s detection ability.
TABLE VI: T-3: New intent discovery result on Test I - OOD set, the best result is in bold, the runner-up result is underlined
Dataset CLINC BANKING SNIPS
[…]
MTP-KNNCL 92.36 79.70 83.60 6min10s/20 83.33 69.75 80.75 6min15s/20 78.04 85.50 96.26 6min40s/20
DPN 89.93 75.19 82.89 6min30s/20 82.51 69.04 76.50 6min/20 68.85 77.72 94.12 6min30s/20
FT-KMeans 86.81 67.71 76.27 4min10s/8 81.88 64.88 73.86 3min/14 66.16 70.29 91.98 3min15s/15
SCL-KMeans (Ours) 93.08 81.14 85.67 1min10s/7 84.44 70.64 81.16 2min30s/10 70.94 75.82 93.58 2min40s/12
V-D Other implementation details
We adopted BERT-base [8] and MPNet-base [57] as the base encoders for both our methods and baselines (expect for PTO-Sup that used GPT-2 small [35]). These encoders were chosen because BERT is the most commonly used pre-trained encoder, while MPNet, a BERT-like variation, enhances context learning ability for better sentence encoding [58]. For SCL setting, we set dropout rate $p\textit{=}0.1$, $n_{\text{views}}\textit{=}2$ for “view” generation, and $\tau\textit{=}0.1$ for calculating the loss. During the pre-training stage, we used the OOD detection AUROC on Val set to evaluate the training process for our SCL and FT methods, saving only the model with the highest $AUROC_{val}$. The maximum training epoch was set to 20 for both our and all baseline methods. All experiments were conducted three times, and the results were averaged to obtain the final result. All experiments were running on the same machine with an Ubuntu 18.04 operating system, an Nvidia A6000 GPU, an Intel i9-10980X CPU, and 128GB RAM.
VI Results and Analysis
VI-A Result on task 1: user intent classification
The intent classification results on Test I-IND data are shown in Table IV. We also provide the training time and epoch as an efficiency reference. It can be found that both proposed SCL and traditional FT method achieved very high F1 scores on CLINC and SNIPS datasets. However, things changed on BANKING, where FT method saw a considerable accuracy drop in both metrics, while our method kept a relatively good performance. We analyze that this is because, compared with the other two datasets, the user inquires in BANKING have a longer input length in average, and the inquiry expressions also become more diverse (that is, more detailed and have many patterns, as the examples shown in Table II), which greatly increased classification difficulty. Despite this, the proposed SCL-NN method was still able to give a more inter-class separated and intra-class compact embedding space (a visual prove will be given in Section VII-A), guaranteeing a high accuracy even with longer and noisier inputs.
VI-B Result on task 2: out-of-domain detection
The results evaluating the detection ability of unknown OOD inputs are shown in Table V. Firstly, comparing FT and SCL, it can be observed that while FT provided very solid results, the proposed method performed even better and achieved nearly all the best results. As for the other baselines, Layer-LOF that applied a layer-voting strategy obtained only slightly better results than FT on the first two datasets. Nevertheless, the cost is a much longer training time by the extensive computations required to measure the information redundancy between layers. The random layer selection strategy also increased uncertainty, making it less practical. On the other hand, the generative model-based method PTO-Sup worked badly except on SNIPS, which even obtained the highest AUROC on it. We speculate the reason of this abnormal situation as compared with the other two datasets, the IND/OOD examples in SNIPS exhibited more “background-shift” than usual “semantic-shift”, which can be exactly the strong point of generation-based methods.
Talking more specifically, for instance, there are two intents belonging to IND and OOD intents shared the topic of “watching a movie” in SNIPS, although one was looking for movie information and another involved planning a cinema visit, which is a typical “background-shift”. (On the contrary, “watching a movie” and “read a book” intent pair is totally different in topic and thus is a “semantic-shift”.) This difference is easier for generative and textual perplexity-based methods like PTO-Sup to capture and distinguish, while it can be a weak point of CE fine-tune-based methods [36]. This is also evident from the performance drop of the other two fine-tuning-based baselines on SNIPS. However, it is clear that SCL still held strong on not only the “semantic-shift”, but also the “background-shift” situation, showing its superiority on the OOD intent detection task.
TABLE VII: T-4: Continual learning result on Test II set
Dataset CLINC BANKING SNIPS
Method/Metric Micro F1$\uparrow$ Macro F1$\uparrow$ Micro F1$\uparrow$ Macro F1$\uparrow$ Micro F1$\uparrow$ Macro F1$\uparrow$
OVERALL
FT 92.65 93.37 79.55 81.69 94.79 94.86
SCL 97.70 97.69 91.75 91.79 99.14 99.14
ON IND
FT 97.53 97.48 86.29 84.33 98.80 98.80
SCL 98.14 98.10 93.53 93.55 99.30 99.29
Ref.-Initial FT model 98.30 98.31 87.11 85.93 99.40 99.40
Ref.-Initial SCL model 98.39 98.40 93.60 93.63 99.40 99.40
ON OOD
FT 89.53 85.65 82.14 78.04 92.18 91.52
SCL 98.48 98.46 92.51 92.30 99.75 99.75
VI-C Result on task 3: new intent discovery
Table VI shows the results for new intent discovery. Comparing SCL and traditional FT method, the advantage of SCL concerning feature optimization capability is proven again by the surpassing clustering scores on all datasets. As for the other baselines, DPN also achieved better scores than FT, but the improvement was not significant. MTP-KNNCL showed very strong clustering performance, achieving second best results to our method on the first two datasets and the best results on SNIPS that other methods performed much worse.
We consider the possible reason as the existence of “background-shift” in SNIPS discussed in T-2 result, where the supervised training style of the other three methods might exacerbate the confusion between those intents. On the other hand, the two techniques, unsupervised MLM target and semi-supervised KNN-ConL, adopted by MTP-KNNCL mitigated the risk of “over-training” in supervised settings, and thus giving a better result. However, the cost is a slightly longer training time.
VI-D Result on task 4: continual learning
Finally, we present the intent classification results after the incremental learning in Table VII. From an overall view (results on all classes), SCL has continued to achieve the best results, and even widened the accuracy gap with FT on all datasets from 0 - 8.5% (on T-1 initial classification) to 4.3% - 10.1%, which we consider as a result of the excellent clustering performance on T-3 and the feature optimization ability of SCL. The concerning forgetting phenomenon also did not appear, as only a slight accuracy decrease on the old IND intent classification was observed after the continual learning. This proves that the adopted “replay” strategy using both pseudo-labeled IND and OOD texts is effective.
In addition to the continuous high accuracy, another advantage of SCL-based re-training is also worth noting. When conducting FT-based re-training, the last fully-connected layer had to be replaced and reinitialized every time to accommodate the new intent number, and it required at least 4-5 epochs to obtain desirable results in our experiments. On the other hand, since SCL focuses on the feature optimization and the label judgment is based on the nearest neighbor algorithm, it can directly use the new data with an extended label set for model re-training in a shorter time. This easy-to-use expandability can save a considerable amount of resources and is meaningful in real-world deployments.
(a) SCL (all test i data)
(b) FT (all test i data)
(c) MTP-KNNCL (all test i data)
(d) SCL (test i - ood data)
(e) FT (test i - ood data)
(f) MTP-KNNCL (test i - ood data)
Figure 4: T-SNE text feature space visualization result on BANKING test I - all data (top row) and only OOD data (bottom row)
VII Additional studies and discussions
VII-A Embedding space visualization
To verify the representation optimization effect, we conducted a t-SNE embedding space visualization on BANKING Test I set, which has a moderate number of examples and categories to view results easily. The results are shown in Figure 4. The top row displays the embedding arrangement over all intents, while the bottom row focuses only on the newly discovered (i.e., previous OOD intents) data.
From the visualization results on overall data distribution (Figure 4 a-c), it can be observed that both SCL and FT performed reasonably well, but the embedding space generated by FT showed more mixed embeddings from different intents. This proves that our proposed method has a better cluster aggregation effect over traditional FT. MTP-KNNCL produced relatively larger clusters in the space with some clusters very close to each other, but the overall distribution situation is slightly better than FT. We consider this may be due to the optimization style of KNN-ConL part used in MTP-KNNCL, which continuously pulled nearby similar embeddings together, resulting in this side effect on cluster aggregation.
From the visualization results of newly discovered intents (Figure 4 d-f), the advantage of proposed method is more obvious. Without accessing OOD texts, SCL was still able to separate most OOD classes, with only two class clusters apparently mixed. MTP-KNNCL also worked relatively well, but the clusters were gathered closer, which showed an inferior labeling dispersion effect to our method. Finally, FT produced a much flatter embedding distribution with more severe confusion between intents. These visualization results correspond exactly to the results in Task 3.
To conclude, the visualization result proved again that despite using less than half the time to the state-of-the-art method MTP-KNNCL and requiring fewer tuning steps, our method still produced a splendid underlying embedding space for clustering and other tasks, which is valuable in practical applications.
VII-B Computational efficiency analysis
Computational efficiency refers to the time or resources required for model training. In this analysis, we use the total computation amount required for a complete training procedure as an indicator of computational efficiency, which can be represented as:
$Eff.\propto Amount\;=\;Epoch\ast(O(model)+O(loss))$ (15)
where O(model) denotes the complexity of model’s forward and back propagation, and O(loss) represents the computation complexity of the loss function.
Since all methods in the experiment used the same base encoder, the O(model) part of them can be considered to be approximately equal [^2] Although GPT-2-small and BERT-base differ, their parameter sizes are close, so we consider their O(model) to be nearly equal.. In addition, given that O(model) is usually significantly larger than O(loss) (The analysis of O(loss) of each method and O(model) can be found in Appendix C), eq. 15 can be simplified to:
$\begin{split}&Eff.\propto Amount\propto Epoch\ast O(model)\\ &\rightarrow Eff.\propto Amount\propto Epoch\end{split}$ (16)
This simplification indicates that the computational efficiency, represented by computation amount, is primarily related to the number of training epochs. Reviewing the results from the tasks (Tables IV-VI), it is evident that the proposed SCL framework consistently converged in fewer epochs and required less training time, and thus demonstrating its advantage on computational efficiency.
Among the baselines, PTO-Sup and Layer-LOF also appeared to be efficient in terms of epoch count, but their training times were significantly longer. We consider the reason as PTO-Sup required complete sentence generation and loss calculation for each word, leading to substantial asynchronous processing delays. Layer-LOF, while not inherently inefficient, used adversarial attacks to oversample the data, which multiplied the training load. These factors inflated their total computation and resulted in suboptimal efficiency.
VII-C Potential challenges, limitations and future solutions
While our method demonstrated strong performance across tasks, as a prototype, it still presents several challenges and limitations needed to be considered for practical deployment.
Background-shift in data: In T-2 and T-3, despite better results than FT, the supervised nature of SCL can still misalign OOD texts with background-shifts, drawing them closer to IND clusters. To address this, future work could include pre-evaluating background-shift intensity and minimizing pre-training epochs to balance IND intent classification and OOD detection accuracy. Incorporating unsupervised training targets may also help.
Noisy inputs: The filtering and re-training strategy between T-3 and T-4 could struggle with noisy inputs that resist categorization, complicating KMeans clustering. The solution could be using advanced clustering methods like HDBSCAN, which can identify outliers without requiring a fixed cluster number. Alternatively, a filtering algorithm could be implemented to iteratively remove distant or scattered data points, though it risks accumulating noises over time.
Scalability on larger datasets: Task 4 results showed good scalability with the addition of up to 38 new intents. However, as category numbers increase, the model’s ability to distinguish between highly similar classes may diminish. When this phenomenon is clearly observed, splitting the model into sub-models based on broader categories may be beneficial, and we plan to explore optimal splitting strategies in future work.
VIII Conclusion
Traditional classification-retrieval-based QA systems primarily focus on categorizing known queries, which is insufficient to address the complexities of real-world scenarios. Moreover, the efficiency of the default training method is often suboptimal. To address these issues, we introduced three new tasks — out-of-domain detection, new intent discovery, and continual learning — to enhance the system’s functional capabilities. We then applied SCL as a unified representation learning method with minimal additional tuning to efficiently achieve the objectives of these tasks. Experimental results demonstrated that, despite its simplicity, the proposed approach consistently outperformed state-of-the-art methods across all tasks. These improvements, integrated within a natural workflow, significantly boost system adaptability and robustness, giving a meaningful impact on the future development of real-world QA systems.
While the proposed method is effective, it remains a prototype with room for further enhancement. Besides the challenges addressed in Section VII-C, future work could focus on integrating more advanced downstream methods to harness the full potential of SCL. For example, a larger-parameter encoder could provide a stronger representation base for learning, while more sophisticated analysis algorithms like energy-based methods should leverage SCL’s tighter cluster formations for enhanced OOD detection effects. In addition, applying advanced techniques like DeepCluster [59] for new intent discovery and Elastic Weight Consolidation [60] for continual learning could also further strengthen the system’s performance. Finally, the proposed SCL framework can also be extended to other QA tasks seamlessly, such as providing top-K confidence results for ambiguous classifications or visualizing sub-intents through embedding space analysis.
Ethical and privacy statement
In this study, the datasets used are all public-available and do not contain privacy-sensitive or confidential information. The neural networks used (i.e., BERT-series) are also less likely to inadvertently disclose sensitive information compared to generative models like GPTs.
Acknowledgment
The authors would like to thank Seita Shimada and Yoshiaki Matsukawa, Rakuten Card Co. Ltd., for their helpful comments on this study.
Appendix A KMeans details
The KMeans learning process is shown in Algorithm 1.
Algorithm 1 KMeans clustering
1:Text embeddings $\{\bm{h}_{j^{\prime}}\}_{j^{\prime}=1}^{N_{ood}}$ of detected $\{\bm{x}_{j^{\prime}}\}_{j^{\prime}=1}^{N_{ood}}$ in T-2, number of clusters $K$
2:Label assignments $\{\hat{y}_{j^{\prime}}\}_{j^{\prime}=1}^{N_{ood}}$
3:Initialize $K$ cluster centroids $\{\bm{\mu}_{k}\}_{k=1}^{K}$ randomly.
4:repeat
5:     for each data point $\bm{h}_{j^{\prime}}$ do
6:         Assign $\bm{h}_{j^{\prime}}$ to the nearest cluster centroid:
$\small\hat{y}_{j^{\prime}}=\arg\min_{k\in\{1,2,\ldots,K\}}\|\bm{h}_{j^{\prime}}-\bm{\mu}_{k}\|^{2}$ (17)
7:     end for
8:     for each cluster $k$ do
9:         Update $\bm{\mu}_{k}$ as the mean of all points assigned to cluster $k$:
$\small\bm{\mu}_{k}=\frac{1}{|num(cate\ k)|}\sum\bm{h}_{j^{\prime}\rightarrow k}$ (18)
10:     end for
11:until the cluster assignments $\{\hat{y}_{j^{\prime}}\}$ do not change or the centroids $\{\bm{\mu}_{k}\}$ converge
12:return Label assignments $\{\hat{y}_{j^{\prime}}\}_{j^{\prime}=1}^{N_{ood}}$
TABLE IX: Abbreviations and full names
Abbreviations Full name
QA question answering (system)
RepL representation learning (feature learning)
ConL contrastive learning
MeL metric learning
UConL unsupervised contrastive learning
SCL supervised contrastive learning
IND in-domain (texts)
OOD out-of-domain (texts)
GDA Gaussian discriminant analysis
LOF local outlier factor
GPT Generative Pre-Trained Transformer
BERT Bidirectional Encoder Representations
from Transformers
CE cross-entropy (loss)
FT fine-tuning
MLM masked language modeling
NN nearest neighbor
Appendix B Metric calculation details
Here we show the detailed calculation processes of the used metrics.
B-A For Task 1 and 4: F1 values
The Micro and Macro F1 scores are calculated as follows:
Micro F1 score: is calculated by summarizing the classification situations over all classes. It is the harmonic mean of micro-averaged precision ($P$) and recall ($R$):
[…]
Macro F1 score: calculates the F1 score independently for each class and then takes the average, which treats all classes equally:
[…]
where $C$ is the number of categories.
B-B For Task 3: NMI, ARI, ACC
The NMI, ARI, ACC are calculated as follows:
Normalized Mutual Information (NMI): is a measure of the similarity between two sets of clusters:
$\text{NMI}(U,V)=\frac{2\cdot I(U;V)}{H(U)+H(V)}$ (21)
where $I(U;V)$ is the mutual information between clusters $U$ and $V$, and $H(U)$ and $H(V)$ are the entropies of $U$ and $V$ respectively. It ranges from 0 (no mutual information) to 1 (perfect correlation).
Adjusted Rand Index (ARI): measures the similarity between two data clusterings by considering all pairs of samples, and counting pairs that are assigned in the same or different clusters in the predicted and true clusterings:
$\text{ARI}=\frac{\sum_{ij}\binom{n_{ij}}{2}-\left[\sum_{i}\binom{a_{i}}{2}\sum_{j}\binom{b_{j}}{2}\right]/\binom{n}{2}}{\frac{1}{2}\left[\sum_{i}\binom{a_{i}}{2}+\sum_{j}\binom{b_{j}}{2}\right]-\left[\sum_{i}\binom{a_{i}}{2}\sum_{j}\binom{b_{j}}{2}\right]/\binom{n}{2}}$ (22)
where $n_{ij}$ is the number of elements in the intersection of clusters $i$ and $j$, $a_{i}$ is the sum over row $i$, and $b_{j}$ is the sum over column $j$.
Clustering Accuracy (ACC): evaluates the clustering result after finding the best one-to-one mapping between cluster labels and true labels:
$\text{ACC}=\frac{\sum_{i=1}^{n}\delta(y_{i},\text{map}(\hat{y}_{i}))}{n}$ (23)
where $y_{i}$ is the true label, $\hat{y}_{i}$ is the cluster label, $\text{map}()$ is the mapping function where we use the Hungary algorithm this time, and $\delta$ is the Kronecker delta function.
Appendix C Computational complexity analysis
The computational complexity analysis of each method is shown in Table VIII according to the original papers.
Here, $N$ is the number of total data, $N_{b}$ is the number of data in one mini-batch and $N_{layer}$ is the number of layers in the network, where for bert-base it is 12. $H$ is the vector dimension of the hidden layer, for bert-base it is 768. $C$ is the number of IND categories. $L$ stands for the average input sentence length, while $V$ is the number of words in dictionary (specially for PTO-Sup). Finally, $K_{adv}$ denotes the number of adversarial examples generated (specially for Layer-LOF), $K_{knn}$ is the number of neighbors used in contrastive learning (specially for MTP-KNNCL) and $K_{proto}$ is the number of assumed prototype categories (specially for DPN).
Appendix D Learning and evaluation procedure (Algorithm form)
The learning and evaluation procedures of pre-training and each task are shown in Algorithm 2.
Appendix E Abbreviations
Table IX shows the abbreviations that have appeared in this paper and their complete names.
Algorithm 2 Overall process
1:Train data, Val data, Test I data, Test II data
2:Solutions for T-1 to T-4
3:      -Pre-training Stage-
4:$Epoch\leftarrow 0$, $AUROC_{best-val}\leftarrow 0$
5:while $Epoch<20$ do
6:     Train ENC on $\bm{x}_{\text{train}}$ using SCL loss object (eq. 8)
7:     Calculate OOD detection $AUROC_{val}$ on this epoch
8:     if $AUROC_{val}>AUROC_{best-val}$ then
9:         $AUROC_{best-val}=AUROC_{val}$
10:         Save model of this epoch, and record the threshold $\lambda$ when TPR=90%
11:     end if
12:     $Epoch=Epoch+1$
13:end while
14:        -Task 1-
15:Input $\bm{x}_{\text{test I}}$ (IND part) to ENC to obtain the text representations:
16:    $\bm{h}_{\text{test I}}=ENC(\bm{x}_{\text{test I}})$
17:The label is determined as the label of the nearest IND cluster centroid:
18:    $\hat{y}_{\text{test I}}=\arg\min_{c\in\mathcal{C}_{kn}}MDist(\bm{h}_{\text{test I}},\bar{\bm{h}}_{c})$
19:Calculate Micro, Macro F1 values
20:        -Task 2-
21:Input $\bm{x}_{\text{test I}}$ (IND, OOD part) to ENC to obtain the representations and calculate the detection scores:
22:    $sco(\bm{x}_{\text{test I}})={\underset{c\in{\mathcal{C}}_{kn}}{min}}\ MDist(\bm{h}_{\text{test I}},\bar{\bm{h}}_{c})$
23:Calculate FPR, TPR results on various thresholds $\lambda$, and finally AUROC, AUPR, FPR90 values
24:      -Task 3 preparation-
25:Determine $\lambda_{filter}$ when OOD detection TPR=90% on $\bm{x}_{\text{val}}$
26:Divide $\bm{x}_{\text{test I}}$ into $\bm{x}_{\text{test I, `IND'}}$ and $\bm{x}_{\text{test I, `OOD'}}$ two parts by comparing detection $sco$ and $\lambda_{filter}$
27:        -Task 3-
28:Apply KMeans on $\bm{x}_{\text{test I, `OOD'}}$ and give new pseudo labels to $\bm{x}_{\text{test I, `OOD'}}$ as $\hat{y}_{\text{test I, `OOD'}}\in\mathcal{C}_{kn^{\prime}}$
29:Calculate NMI, ARI, ACC metrics
30:        -Task 4-
31:Give pseudo labels to $\bm{x}_{\text{test I, `IND'}}$ using the initial ENC as $\hat{y}_{\text{test I, `IND'}}\in\mathcal{C}_{kn}$
[…]
33:Test on $\bm{x}_{\text{test II}}$ and calculate Micro, Macro F1 values
34:END
[…]